OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1

OpenStudy (diyadiya):

Add the two equation 3x + y = 14 –3x + 2y = 1 ------------ 0 +3y = 15 3y =15 Divide both sides by 5 y= ?

OpenStudy (diyadiya):

@reggie21 Did you get the value of y ?